OpenStudy (anonymous):

1. If a and b are any two real numbers, explain how you calculate the difference a – b. Please answer all parts to the problem. a. How are addition and subtraction related? b. How you can determine the sign of your solution? c. How you can use a number line to help with the calculation? Include an example using a value for both a and b.

jimthompson5910 (jim_thompson5910):

Rules: Rule 1) if both a & b are positive, and a > b, then you just subtract like normal if both a & b are positive, and a < b, then you just subtract (larger - smaller) and you make the result negative ------------------- Rule 2) if a is positive and b is negative, then a-b will turn into an addition problem example: let a = 2, b = -3 a-b = 2 - (-3) = 2+3 = 5 ------------------- Rule 3) if a is negative and b is positive, then you add up the terms without the negative sign and then make the result negative example: let a = -7 and b = 10 a-b = -7 - 10 = -17 notice how i did 7+10 = 17 and made the result negative -------------------- Rule 4) if a & b are both negative then you'll refer to rule 1 example a = -2, b = -5 a - b = -2 - (-5) = -2 + 5 = 5 - 2 = 3